OpenStudy (anonymous):

The negatives are confusing me in this problem. Write the equation of the quadratic function with roots -1 and -7 and a vertex at (-4, 7)

OpenStudy (phi):

roots -1 and -7 that means x= -1 and x= -7 you can rewrite x= -1 by adding +1 to both sides . you get x+1 = -1+1 or x+1= 0 and x= -7 means (x+7)=0 if we multiply (x+1)(x+7) we know this is still equal to 0 (x+1)(x+7)=0 if we multiply by some number, (call it a), we get a(x+1)(x+7)= a*0 or 0= a(x+1)(x+7) to make a quadratic equation, let this be equal to y y= a(x+1)(x+7) the last step is find a. use the vertex info (-4, 7), which means when x=-4, y = 7 can you do that ?

OpenStudy (anonymous):

Um, 7= a(-4 + 1)(-4+7)?

OpenStudy (phi):

as a first step. simplify

OpenStudy (anonymous):

simplify 7= a(-4 + 1)(-4+7)?

OpenStudy (phi):

yes.

OpenStudy (anonymous):

7=a(-3)(3) 7=a(-9)?

OpenStudy (phi):

one more step. divide both sides by -9

OpenStudy (anonymous):

a=-7/9

OpenStudy (phi):

so one form of the equation is y= a(x+1)(x+7) with a= -7/9 y = (-7/9) (x+1)(x+7) if this is multiple choice it may not be easy to match this to your choices, if the choices are in a different form.

OpenStudy (phi):

The other way to get the equation is use the vertex form : y= a(x-h)^2 + k where (h,k) is the vertex you replace h= -4, and k= 7 y = a(x- (-4))^2 + 7 or y= a(x+4)^2 + 7 a will be the same as before, but if we did not know it. try the point (-1,0) (one of the roots)

OpenStudy (phi):

for example, \[ y= a(x+4)^2 + 7 \] use (-1,0) to get \[ 0 = a(-1+4)^2 + 7\\ 0= 9a + 7\\ 9a= -7\\ a= -\frac{7}{9} \]

OpenStudy (anonymous):

It is a written question buy that is not the form we are supposed to write the final equation in..

OpenStudy (phi):

what form do they want ?

OpenStudy (anonymous):

well an example is roots being 6 and 8 and vertex (7, -5) The final equation is y=5x^2-70x+240

OpenStudy (anonymous):

The hardest part for me is writing the final equation correctly, I have done a ton of these problems.

OpenStudy (phi):

using vertex form for the problem well an example is roots being 6 and 8 and vertex (7, -5) y = a( x - 7)^2 -5 use (6,0) to find a: 0= a(6-7)^2 -5 0 = a* (-1)^2 -5 5= 1a a=5 y= 5(x-7)^2 -5 to change from vertex form to standard form expand the (x-7)^2= x^2 -14x +49 y= 5(x^2 -14x +49)-5 y = 5x^2 -70 x +(5*49-5) y = 5x^2 -70 x + 240

OpenStudy (anonymous):

How would I come up with the final equation for the other problem in this form?

OpenStudy (phi):

\[ y= - \frac{7}{9} \left(x+4\right)^2 + 7 \] expand (x+4)(x+4)= x^2 +8x +16 \[ y= - \frac{7}{9} \left(x^2 +8x +16\right) + 7 \] I would make the final 7 = 63/9 so we can write this as \[ y= - \frac{7}{9} \left(x^2 +8x +16\right) + \frac{63}{9}\\ y= \frac{1}{9}\left(-7x^2 -56x-112 +63\right)\] \[y= \frac{-7x^2 -56x-49}{9}\]

OpenStudy (anonymous):

Thank You so much!